OpenStudy (anonymous):

Help I am having trouble on my homework for algebra.!!! I will give a medal and a fan if you can help me:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

$100 at 6.5% for 3 years

OpenStudy (anonymous):

how am i suposed to answer these

OpenStudy (anonymous):

are you in chapter 4

OpenStudy (anonymous):

can u type the whole question, i am not clear with the question

OpenStudy (anonymous):

it says to calculate the simplest interest

OpenStudy (anonymous):

are you in chapter 4

OpenStudy (anonymous):

I have no idea what chapter

OpenStudy (anonymous):

I have a few more after this that I will re-post so you guys can get more medals :)

OpenStudy (anonymous):

105.0 i think

OpenStudy (anonymous):

I THINK

OpenStudy (anonymous):

are you 100% sure this is the last chance I get for this question I got it wrong 2 times already.

OpenStudy (anonymous):

use the formula I=pnr /100 where p=amount n=no. of years r=%ge

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I=100*6.5%*3?????????

OpenStudy (anonymous):

sorry but i dont know then just ask samb

OpenStudy (anonymous):

I =interest and substitute the known values

OpenStudy (anonymous):

@kyley_bear the I=(100*6.5*3)/100

OpenStudy (anonymous):

I=6.5%(100*3)

OpenStudy (anonymous):

hence I = 19.5

OpenStudy (anonymous):

be sure u remember the formula @kyley_bear

OpenStudy (anonymous):

OHHHH okay :)

OpenStudy (anonymous):

THERE GO YOUR MEDAL!!!!!!!

OpenStudy (anonymous):

thanks okay Ill go post another :')

OpenStudy (anonymous):

CAN U BE MY FAN

OpenStudy (anonymous):

okay :) I fanned you both :)